# Simulation-based inference of galaxy properties from JWST pixels

### v1

@@ 6.1 Maps of stellar population properties
Figure 5: Galaxy property maps. We include only the pixels with S/N (averaged between F277W, F356W and F [^4] W) higher than 5. We use a $\tau$-delayed prior for the SFHs. From left to right we show the RGB images (built using the method described in Lupton et al. (2004) with the filters specified above each image to enhance the different structures), and the inferred properties per pixel $\log_{10}\left(\Sigma_{\star}/\left(M_{\odot}\mathrm{kpc}^{-2}\right)\right)$, the mass-weighted age [Gyr], and $A_{V}$ [mag]. In the lower-left corner of the first column, we show circles with diameters equal to the FWHM of the F [^4] W PSF. Scale bars in the first column indicate angular (arcseconds) and physical (kiloparsecs) sizes.
@@
Figure 6: Same as Fig. 5 but for the Dirichlet prior for the SFHs. We infer the properties $\log_{10}\left(\Sigma_{\star}/\left(M_{\odot}\mathrm{kpc}^{-2}\right)\right)$, $t_{50\%}$ [Gyr] in lookback time, and $A_{V}$.
[…]
@@
In contrast, for galaxies at higher redshifts (117960, 118081 and 206146), our analysis primarily captures the effective radii, which is of the order of the PSF FWHM for F [^4] W ($0.16^{\prime\prime}$, $5.36$ pix), because of the S/N threshold. Consequently, the maps reveal smooth trends, with finer structural gradients remaining unresolved. The mass density maps show less gradients. The galaxies 117960 and 118081 present signs of inside-ouside star formation, with cores younger than the outskirts. The dust attenuation is less strong for these galaxies, with smaller variations and slight negative correlations with the age, and with younger regions more attenuated. Additionally, we show in Appendix D the maps of uncertainties for the properties of the six galaxies that we obtain from the posterior distributions.
@@
Continuing with the study of the six galaxies, we show in Fig. 6 the maps of the stellar population properties obtained with the model using the Dirichlet prior for the SFHs. The structures we find in these maps are very similar to those obtained with the $\tau$ delayed prior. However, we find more structure in the stellar mass maps for the high-redshift galaxies. The other differences are mainly in the age, which is expected from the differences in the SFH shapes, which in this cased is computed as $t_{50\%}$, the lookback time in Gyr at which 50% of the total stellar mass was formed, itself a direct parameter of the model.
@@
We do not observe significant variations within the PSF size of the F [^4] W filter, and all variations are consistent with photometric noise levels, demonstrating the robustness of the inference. Although this behaviour was not explicitly enforced in our model, as is done in some binned modelling approaches, it naturally arises from our methodology. These maps offer multiple pathways for interpreting pixel-by-pixel properties within the global context of each galaxy. For example, rather than binning the photometry, we could directly bin the inferred properties, leveraging the strengths of Bayesian inference and hierarchical modelling to narrow the posterior distributions without losing the high resolution of our photometry. Alternatively, we could analyse gradients or property profiles to explore how these quantities vary across the galaxy. As a proof of concept, we include in Appendix F the radial profiles of the properties shown in the maps.
@@ 6.2 Comparison between resolved and integrated stellar masses
Figure 7: From top to bottom, for $1083$ galaxies, the pixel-based vs the integrated estimations for the stellar masses for the $\tau$-delayed prior for the SFHs, for the Dirichlet prior, the pixel-based stellar masses for the two different priors and the integrated stellar masses for the two priors. Each galaxy is a different dot colour-coded with the redshift. We also show the six previously studied galaxies as stars, as well as black KDE contours and in red the 1:1 ratio (solid line) and ±$\pm 0.5$ dex (dashed lines) and ±$\pm 1$ dex (dotted lines) offsets. For clarity, only the error bars of the six galaxies are shown for both the integrated and resolved stellar masses.
@@
We now study all the pixels of the $1083$ galaxies with $\overline{\rm{S/N}}$¿5, $\sim 2$ million pixels. We focus on the surviving stellar mass recovery, leaving the analysis of other properties for future exploration. Our approach treats individual pixels as statistically independent for simplicity and computational feasibility. While we recognize that the PSF inherently introduces spatial correlations between neighbouring pixels, we will thoroughly investigate this in upcoming works. For now, we compute the total stellar mass by summing posterior samples across pixels.
@@
Specifically, we draw one sample from the posterior distribution of each pixel and sum these to obtain one sample of the posterior for the galaxy’s total stellar mass. Repeating this process for all posterior samples yields a posterior distribution for the total mass. We then use the medians of this total-mass posteriors to compare with the integrated fit. For the latter, we sum the photometry of each of the above pixels and compute the integrated stellar masses for all the galaxies. We do this for the two models with different priors for the SFHs to compare the differences between these two ways of computing total stellar masses.
@@
Several works attribute systematic differences between those stellar masses due to outshining of young stellar populations in the integrated analysis (Sorba & Sawicki, 2015, 2018; Jain et al., 2024; Giménez-Arteaga et al., 2024; Fujimoto et al., 2024; Narayanan et al., 2024), as a consequence of bright O-type and B-type stars that obscure the old populations which dominate the total stellar mass. This outshining phenomenon is thought to be stronger in high-redshift low-mass galaxies, due to the stochasticity in their SFHs. Other recent works claim that this problem is mainly solved due to the large wavelength coverage of JWST (Song et al., 2023; Lines et al., 2025) and medium bands of NIRCam (Adams et al., 2025; Cochrane et al., 2025; Harvey et al., 2025). With our approach we can perform such a study in a Bayesian framework for a statistically significant set of observed galaxies, with at least one order of magnitude more galaxies than previously.
@@
We show in Fig. 7 the resolved vs integrated stellar masses for all galaxies colour-coded with redshift, as well as the six galaxies studied in detail before with symbols of stars, for both the $\tau$-delayed and the Dirichlet prior for the SFHs. We do not find strong outshining effects for any of the priors for stellar masses from $10^{8}M_{\odot}$ to $10^{12}M_{\odot}$. However, the scatter is larger for the less flexible $\tau$-delayed SFHs. For this prior we find $10$ outliers with offset (computed as $\log_{10}\left(M_{*}^{\text{resolved}}\right)-\log_{10}\left(M_{*}^{\text{integrated}}\right)$) higher than $1$ dex (one outlier with $z<3$ and nine with $z>3$). The median offset for this prior is $-0.068$ dex for $z<3$ and $0.063$ dex for $z>3$, while the median standard deviation for the posterior distributions of integrated stellar mass are respectively $0.14$ dex and $0.20$ dex. The uncertainties associated to the resolved masses, derived from the uncertainty propagation of the standard deviation of the posterior distributions of each of the pixels, are instead $0.044$ dex ($z<3$) and $0.14$ dex ($z>3$). For the Dirichlet prior, we find five outliers, all of them with $z>3$. The median offsets are $0.091$ dex for $z<3$ and $0.016$ for $z>3$, while the median standard deviation for the posterior distributions of integrated stellar mass are respectively $0.20$ dex and $0.31$ dex, and the uncertainties of the resolved analysis are $0.014$ dex and $0.089$ dex (from the uncertainty propagation of the standard deviations of the posteriors). The largest offsets and all the outliers are found for low-mass galaxies, below $10^{8}M_{\odot}$. This is consistent with the findings of Lines et al. (2025) and Harvey et al. (2025).
@@
Fig. 7 also compares the resolved and integrated stellar masses estimated for the same galaxies using the two different SFH priors. We measure a significant offset between the resolved stellar masses estimated with a $\tau$-delayed and a Dirichlet prior. The median offsets are of $0.204$ dex at $z<3$ and of $0.028$ dex at $z>3$. The differences are less pronounced when comparing stellar masses obtained from integrated photometry. We report median offsets of $0.044$ dex at $z<3$ and $-0.072$ dex at $z>3$.
@@
To further investigate this we perform a cross-test in the simulations, that we describe in Appendix 16. The results in the simulations show that, due to the flexibility of the model trained on the Dirichlet simulation, it can correctly estimate the masses from a simulation performed with a $\tau$-delayed prior, marginalising over all parameters that the model cannot constrain from photometry (such as possible SFHs) and accurately reflecting this in the uncertainties. In contrast, the model using the $\tau$-delayed prior systematically underestimates the stellar masses when applied to the Dirichlet simulation, on average by $0.147$ dex, roughly consistent with the reported offsets in observations. By inspecting the inferred and true SFHs, we conclude that the $\tau$-delayed prior leads to model misspecification: it only recovers the last burst of star formation and tends to ignore significant fractions of stellar mass formed in early epochs, without accounting for them in the uncertainties of the recovered properties.
7 Summary, limitations and future work
We demonstrate the efficiency and robustness of simulation-based inference for recovering the properties of simulated and observed stellar populations from multi-wavelength photometric data. A key strength of our approach lies in its computational efficiency. The amortized inference enabled by NFs allows for the rapid analysis of vast datasets. We achieved a fitting time of approximately $0.04$ seconds per pixel, a speed that is $\sim 3.5\times 10^{3}$ times faster than traditional Bayesian codes like Prospector. This speed is crucial for the pixel-by-pixel analysis of thousands of galaxies, enabling a statistical Bayesian approach to resolved stellar population studies that would be prohibitively time-consuming with traditional methods.
@@ 7 Summary, limitations and future work
The amortized nature of our methodology, while offering a significant computational advantage over traditional Bayesian approaches like MCMC, imposes constraints on its applicability. The model’s training is the only computationally intensive step, enabling inference speeds orders of magnitude faster. However, it requires retraining for different filter sets or observational setups, as the simulation must precisely match the observed data, including uncertainties and background noise. Expanding the simulation to encompass multiple surveys would significantly increase the computational burden of training. Nonetheless, the substantial sample sizes available within surveys like JADES DR2 (approximately $94,000$ objects, with $\sim 11,000$ meeting our filter requirements) justify this approach, marking a notable advancement. We plan to extend this pipeline to other JWST surveys, such as The Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al., 2025), and explore fields with different observational properties.
Applying our model to 1083 galaxies with spectroscopic redshifts, we explore the properties per pixel and their dependence on S/N at different redshifts. Stellar masses, star formation rates, mass-weighted ages and dust attenuation can be well constrained down to $\overline{\rm{S/N}}=5$, while the recovery of other properties, such as metallicity, is more challenging. The choice of SFH prior impacts the inferred properties, including the stellar mass, which shows an average offset of 0.15 dex in our cross-validation study. This highlights the importance of considering prior assumptions when assessing uncertainties.
@@
A primary limitation in our galaxy sample selection was the reliance is spectroscopic redshifts, which are sparsely available. While this choice minimises redshift uncertainty in SED fitting, it restricts the number of galaxies analysed. To address this, one could incorporate photometric redshifts, either as direct inputs without retraining the model, or retraining it with their associated uncertainties as input of the network. However, this will introduce a larger uncertainty into other property estimations. Furthermore, our assumption of equal redshift across all pixels within a galaxy introduces another source of potential error. While we generally trust the segmentation maps, galaxy overlaps or ambiguous pixel assignments can occur. Inferring redshifts per pixel or clump, as demonstrated in Pérez-González et al. (2023), offers a potential solution. Although this approach may be susceptible to low S/N issues, we intend to investigate it in future works, seeking to detect and resolve redshift discrepancies indicative of contamination or to refine photometric redshift estimates through Bayesian combination of posterior distributions.
We show-case the model’s ability to resolve spatial variations in galaxy physical properties through pixel-by-pixel analysis of six example JADES galaxies across different redshifts and morphologies. We observe clear gradients in stellar mass, age, and dust attenuation, particularly in the low-redshift galaxies $(z<2)$ where the PSF FWHM for the F [^4] W filter (to which all the photometry mosaics were convolved) is considerably smaller than the galaxy size.
As a proof of concept for future statistical analysis, we compare pixel-by-pixel and integrated estimations of stellar masses for the $1083$ galaxies in our sample and their $\sim 2$ million pixels with $\overline{\rm{S/N}}>5$. We find no strong evidence of outshining effects, with only a few outliers observed in high-redshift low-mass galaxies. This result is consistent with recent studies that have leveraged the high wavelength coverage of HST+JWST including NIRCam’s medium bands (e.g., Song et al., 2023; Lines et al., 2025; Cochrane et al., 2025; Harvey et al., 2025). However, it deviates from other findings that suggested significant outshining (Sorba & Sawicki, 2015, 2018; Jain et al., 2024; Giménez-Arteaga et al., 2024; Narayanan et al., 2024). We attribute the primary systematic uncertainty in stellar mass estimation to the choice of SFH prior, rather than the resolved or integrated analysis.
@@
The choice of SFH prior, whether restrictive (e.g., $\tau$-delayed) or flexible (e.g., Dirichlet), influences the retrieved properties and their uncertainties. While this study explored these extremes, real galaxy SFHs may exhibit intermediate complexity. More flexible priors, while potentially more realistic, often lead to wider posterior distributions and increased uncertainty. In the recent results of Harvey et al. (2025), the continuity and power-law priors were the ones that lead to less outshining effects. We aim to explore these priors in future works. Furthermore, the choice of SPS model can significantly influence results, as seen in comparisons with Synthesizer. These limitations inherent in SED models reflect our incomplete understanding of stellar evolution, a factor that equally affects conventional MCMC sampling approaches and extends beyond the scope of this work.
Finally, we address the necessity of pixel-by-pixel analysis. For NIRCam the PSF size in all bands is significantly larger than the pixel scale ($0.03^{\prime\prime}$), leading to potential misinterpretations at individual pixel scales. While we do not find variations under the PSF FWHM scales, our model’s adaptability to integrated or binned data without retraining is a key advantage. Beyond traditional binning methods like Voronoi tessellation (Cappellari & Copin, 2003), which fails to exploit the high resolution of the images by ensuring a constant S/N, the piXedfit pixel-binning methodology (Abdurro’uf & Akiyama, 2017; Abdurro’uf et al., 2021) offers a promising alternative. By segmenting galaxies into components based on $\chi^{2}$ comparisons of neighbouring SEDs, it maximizes spatial SED information while respecting PSF limitations and maintaining sufficient S/N within each bin. We plan to explore this approach in future investigations.

### v3

@@ 6.1 Maps of stellar population properties
Figure 5: Galaxy property maps. We include only the pixels with a S/N (averaged between F277W, F356W, and F [^4] W) higher than 5. We use a $\tau$-delayed prior for the SFHs. From left to right: RGB images (built using the method described in Lupton et al. (2004) with the filters specified above each image to enhance the different structures), and the inferred properties per pixel $\log_{10}\left(\Sigma_{\star}/\left(M_{\odot}\mathrm{kpc}^{-2}\right)\right)$, the mass-weighted age [Gyr], and $A_{V}$ [mag]. In the lower-left corner of the first column, we show circles with diameters equal to the FWHM of the F [^4] W PSF. Scale bars in the first column indicate angular (arcseconds) and physical (kiloparsecs) sizes.
@@
Figure 6: Same as Fig. 5, but for the Dirichlet prior for the SFHs. We infer the properties $\log_{10}\left(\Sigma_{\star}/\left(M_{\odot}\mathrm{kpc}^{-2}\right)\right)$, $t_{50\%}$ [Gyr] in lookback time, and $A_{V}$.
After comparing our model with Prospector and Synthesizer for a few pixels, we proceeded to use it to create maps of the stellar population properties of the six sample galaxies. We also built RGB images using a colour scheme designed to enhance the structures enabling a better comparison with the population maps. We show the RGB images and the medians of the posteriors retrieved for the three physical properties for the six galaxies in Fig. 5, using the parametric SFHs. For the galaxies at lower redshifts (254985, 205449 and 211273), the PSF full width at half maximum (FWHM) is considerably smaller than the visual size of the galaxies. This, when combined with the large number of pixels available, enables clear gradients to be resolved in all the properties. The stellar mass density, $\log_{10}\left(\Sigma_{\star}/\left(M_{\odot}\mathrm{kpc}^{-2}\right)\right)$, is concentrated predominantly in the central regions of galaxies. The ages of the stellar populations are not as symmetric and are, instead, aligned with the morphology of the galaxies, with young populations along the spiral arms for the first two galaxies, which are face-on, and in the disc plane for the more inclined third galaxy. The dust attenuation shows strong variations and is correlated with the age, with young dusty clumps, together with centres and inner spiral arms highly attenuated with values up to $\penalty 10000\ 4$ mag for 211273.
@@
In contrast, for galaxies at higher redshifts (117960, 118081, and 206146), our analysis primarily captures the effective radii, which is of the order of the PSF FWHM for F [^4] W ($0.16^{\prime\prime}$, $5.36$ pix) because of the S/N threshold. Consequently, the maps reveal smooth trends, with the finer structural gradients remaining unresolved. The mass density maps show fewer gradients. The galaxies 117960 and 118081 present signs of inside-outside star formation, with cores younger than the outskirts. The dust attenuation is less strong for these galaxies, with smaller variations and slight negative correlations with the age, while for younger regions, it is more attenuated. Additionally, we show in Appendix D the maps of uncertainties for the properties of the six galaxies that we obtained from the posterior distributions.
@@
Continuing with the study of the six galaxies, we show in Fig. 6 the maps of the stellar population properties obtained with the model using the Dirichlet prior for the SFHs. The structures we find in these maps are very similar to those obtained with the $\tau$ delayed prior. However, we find more structure in the stellar mass maps for the high-redshift galaxies. The other differences are mainly in terms of the age, which is expected from the differences in the SFH shapes; in this case, this is computed as $t_{50\%}$, which is the lookback time in Gyr at which 50% of the total stellar mass was formed (itself a direct parameter of the model).
@@
We did not observe significant variations within the PSF size of the F [^4] W filter and all the variations are consistent with photometric noise levels, demonstrating the robustness of the inference. Although this behaviour was not explicitly enforced in our model (as is sometimes done in some binned modelling approaches), it naturally arises from our methodology. These maps offer multiple pathways for interpreting pixel-by-pixel properties within the global context of each galaxy. For example, rather than binning the photometry, we could directly bin the inferred properties, leveraging the strengths of Bayesian inference and hierarchical modelling to narrow the posterior distributions without losing the high resolution of our photometry. Alternatively, we could analyse gradients or property profiles to explore how these quantities vary across the galaxy. As a proof of concept, we include the radial profiles of the properties shown in the maps in Appendix F.
@@ 6.2 Comparison between resolved and integrated stellar masses
Figure 7: Pixel-based vs the integrated estimations for the stellar masses for the $\tau$-delayed prior for the SFHs, for the Dirichlet prior, the pixel-based stellar masses for the two different priors, and the integrated stellar masses for the two priors for $1083$ galaxies (from top to bottom). Each galaxy is a different dot colour-coded with the redshift. We also show the six previously studied galaxies as stars, as well as black kernel density distribution contours, and in red the 1:1 ratio (solid line), $\pm 0.5$ dex (dashed lines), and $\pm 1$ dex (dotted lines) offsets. For clarity, only the error bars of the six galaxies are shown for both the integrated and resolved stellar masses.
@@
We went on to study all the pixels of the $1083$ galaxies with $\overline{\rm{S/N}}$¿5, $\sim 2$ million pixels. We focussed on the surviving stellar mass recovery, leaving the analysis of other properties for future exploration. Our approach treats individual pixels as statistically independent for simplicity and computational feasibility. While we recognise that the PSF inherently introduces spatial correlations between neighbouring pixels, we plan to thoroughly investigate this in upcoming works. For the purposes of this work, we chose to compute the total stellar mass by summing the posterior samples across pixels.
@@
Specifically, we drew one sample from the posterior distribution of each pixel and summed these to obtain one sample of the posterior for the galaxy’s total stellar mass. Repeating this process for all posterior samples yields a posterior distribution for the total mass. We then used the medians of this total-mass posteriors to compare them with the integrated fit. For the latter, we summed the photometry of each of the above pixels and computed the integrated stellar masses for all the galaxies. We did this for the two models with different priors for the SFHs to compare the differences between these two ways of computing total stellar masses.
@@
Several works attribute systematic differences between those stellar masses due to outshining of young stellar populations in the integrated analysis (Sorba & Sawicki, 2015, 2018; Jain et al., 2024; Giménez-Arteaga et al., 2024; Fujimoto et al., 2025; Narayanan et al., 2024), as a consequence of bright O-type and B-type stars that obscure the old populations, which dominate the total stellar mass. This outshining phenomenon is thought to be stronger in high-redshift low-mass galaxies, due to the stochasticity in their SFHs. Other recent works claim that this problem is mainly solved due to the large wavelength coverage of JWST (Song et al., 2023; Lines et al., 2025) and medium bands of NIRCam (Adams et al., 2025; Cochrane et al., 2025; Harvey et al., 2025). Our approach allowed us to perform such a study in a Bayesian framework for a statistically significant set of observed galaxies, comprising at least one order of magnitude more galaxies than previous works.
@@
We show in Fig. 7 the resolved versus integrated stellar masses for all galaxies colour-coded with redshift, as well as the six galaxies studied in detail before with symbols of stars, for both the $\tau$-delayed and the Dirichlet prior for the SFHs. We did not find any strong outshining effects for any of the priors for stellar masses from $10^{8}M_{\odot}$ to $10^{12}M_{\odot}$. However, the scatter is larger for the less flexible $\tau$-delayed SFHs. For this prior, we found ten outliers with their offsets (computed as $\log_{10}\left(M_{*}^{\text{resolved}}\right)-\log_{10}\left(M_{*}^{\text{integrated}}\right)$) higher than $1$ dex (one outlier with $z<3$ and nine with $z>3$). The median offset for this prior is 0.068-0.068 dex for $z<3$ and $0.063$ dex for $z>3$, while the median standard deviation for the posterior distributions of integrated stellar mass are $0.14$ dex and $0.20$ dex, respectively. The uncertainties associated to the resolved masses, derived from the uncertainty propagation of the standard deviation of the posterior distributions of each of the pixels, are instead $0.044$ dex ($z<3$) and $0.14$ dex ($z>3$). For the Dirichlet prior, we found five outliers, all of them with $z>3$. The median offsets are $0.091$ dex for $z<3$ and $0.016$ for $z>3$, while the median standard deviation for the posterior distributions of integrated stellar mass are $0.20$ dex and $0.31$ dex, respectively, and the uncertainties of the resolved analysis are $0.014$ dex and $0.089$ dex (from the uncertainty propagation of the standard deviations of the posteriors). The largest positive offsets and all the outliers were found for low-mass galaxies (below $10^{8.5}M_{\odot}$) at all redshifts, reporting offsets of 0.40 dex and 0.25 dex for the parametric and non-parametric models, respectively; this can be compared with the offsets for larger stellar masses, namely, of 0.11 dex and 0.09 dex. This is consistent with the findings of Lines et al. (2025) and Harvey et al. (2025).
@@
In Fig. 7, we also compare the resolved and integrated stellar masses estimated for the same galaxies using the two different SFH priors. We measured a significant offset between the resolved stellar masses estimated with a $\tau$-delayed and a Dirichlet prior. The median offsets are of $0.204$ dex at $z<3$ and of $0.028$ dex at $z>3$. The differences are less pronounced when comparing stellar masses obtained from integrated photometry. We reported median offsets of $0.044$ dex at $z<3$ and 0.072-0.072 dex at $z>3$.
@@
To further investigated this, we performed a cross-test in the simulations, which we describe in Appendix 16. The results in the simulations show that due to the flexibility of the model trained on the Dirichlet simulation, it can correctly estimate the masses from a simulation performed with a $\tau$-delayed prior, marginalising over all parameters that the model cannot constrain from photometry (e.g. possible SFHs) and accurately reflecting this in the uncertainties. In contrast, the model using the $\tau$-delayed prior systematically underestimates the stellar masses when applied to the Dirichlet simulation, on average, by $0.147$ dex; this is roughly consistent with the reported offsets in observations. By inspecting the inferred and true SFHs, we conclude that the $\tau$-delayed prior leads to model misspecification: it only recovers the last burst of star formation and tends to ignore significant fractions of stellar mass formed in early epochs, without accounting for them in the uncertainties of the recovered properties.
7 Limitations
@@
The amortised nature of our methodology, while offering a significant computational advantage over traditional Bayesian approaches such as MCMC, imposes constraints on its applicability. The model’s training is the only computationally intensive step, enabling inference speeds orders of magnitude faster. However, it requires retraining the model for different filter sets or observational setups, as the simulation must precisely match the observed data, including uncertainties and background noise. Expanding the simulation to encompass multiple surveys would significantly increase the computational burden of training. Nonetheless, the substantial sample sizes available within surveys such as JADES DR2 (approximately $94,000$ objects, with $\sim 11,000$ meeting our filter requirements) justify this approach, marking a notable advancement.
@@
A primary limitation in our galaxy sample selection was the reliance on spectroscopic redshifts, which are sparsely available. While this choice minimises redshift uncertainty in SED fitting, it restricts the number of galaxies analysed. To address this, we could incorporate photometric redshifts, either as direct inputs without retraining the model or retraining it with their associated uncertainties as input of the network. However, this will introduce a larger uncertainty into other property estimations. Furthermore, our assumption of equal redshift across all pixels within a galaxy introduces another source of potential error. While we generally trust the segmentation maps, galaxy overlaps or ambiguous pixel assignments can occur. Inferring redshifts per pixel or clump, as demonstrated in Pérez-González et al. (2023), offers a potential solution. Although this approach may be susceptible to issues of low S/N, we intend to investigate it in future works, seeking to detect and resolve redshift discrepancies indicative of contamination or to refine photometric redshift estimates through Bayesian combination of posterior distributions.
@@
The choice of SFH prior, whether restrictive (e.g. $\tau$-delayed) or flexible (e.g. Dirichlet), influences the retrieved properties and their uncertainties. While this study explored these extremes, real galaxy SFHs may exhibit intermediate complexity. More flexible priors, while potentially more realistic, often lead to wider posterior distributions and increased uncertainty. In the recent results of Harvey et al. (2025), the continuity and power-law priors were the ones that lead to less outshining effects. We aim to explore these priors in future works. Furthermore, the choice of SPS model can significantly influence results, as seen in comparisons with Synthesizer. These limitations inherent in SED models reflect our incomplete understanding of stellar evolution, a factor that equally affects conventional MCMC sampling approaches and extends beyond the scope of this work.
An important consideration to take into account is whether fitting individual pixels without modelling the PSF introduces biases. Our framework does not require explicit PSF characterisation in the forward model because each pixel is treated as an independent stellar population, rather than as part of a resolved spatial structure. However, oversampling the PSF can introduce statistical dependencies between neighbouring pixels. This does not bias individual posterior estimates but could affect summed quantities, such as total stellar mass or SFRs, due to correlated errors or partial double-counting of flux, especially in compact sources. To quantify this effect, we would need cosmological hydrodynamical simulations that properly model galaxy morphology, physical properties, and observational effects, allowing us to compare pixel-by-pixel and integrated estimates against known ground truth. We could also address this issue with appropriate spatial binning or post-processing to account for PSF scale correlations. The SBI framework allows for posterior-based binning after inference, for instance, using measures such as the Kullback-Leibler divergence between posteriors rather than simple spatial proximity, enabling more adaptive grouping strategies that respect both the S/N level and physical continuity. We plan to explore these two approaches in future investigations.
8 Summary and future work
We demonstrate the efficiency and robustness of simulation-based inference for recovering the properties of simulated and observed stellar populations from multi-wavelength photometric data. A key strength of our approach lies in its computational efficiency. The amortised inference enabled by NFs makes it possible to carry out rapid analyses of vast datasets. We achieved a fitting time of approximately $0.04$ seconds per pixel, which is a speed that is $\sim 3.5\times 10^{3}$ times faster than traditional Bayesian codes such as Prospector. This speed is crucial for the pixel-by-pixel analysis of thousands of galaxies, enabling a statistical Bayesian approach to resolved stellar population studies that would be prohibitively time-consuming with traditional methods.
Applying our model to 1083 galaxies with spectroscopic redshifts, we explored the properties per pixel and their dependence on S/N at different redshifts. Stellar masses, SFRs, mass-weighted ages, and dust attenuation can be well constrained down to $\overline{\rm{S/N}}=5$, while the recovery of other properties, such as metallicity, is more challenging. The choice of SFH prior impacts the inferred properties, including the stellar mass, which shows an average offset of 0.15 dex in our cross-validation study. This highlights the importance of considering prior assumptions when assessing the uncertainties.
We showcased the model’s ability to resolve spatial variations in galaxy physical properties via a pixel-by-pixel analysis of six example JADES galaxies across different redshifts and morphologies. We observed clear gradients in terms of stellar mass, age, and dust attenuation, particularly in the case of low-redshift galaxies $(z<2),$ where the PSF FWHM for the F [^4] W filter (to which all the photometry mosaics were convolved) is considerably smaller than the galaxy size.
As a proof of concept for future statistical analysis, we made a pixel-by-pixel comparison and integrated the estimations of stellar masses for the $1083$ galaxies in our sample and their $\sim 2$ million pixels with $\overline{\rm{S/N}}>5$. We found no strong evidence of outshining effects, with only a few outliers observed in high-redshift low-mass galaxies. This result is consistent with recent studies that have leveraged the high-wavelength coverage of HST+JWST, including NIRCam’s medium bands (e.g. Song et al., 2023; Lines et al., 2025; Cochrane et al., 2025; Harvey et al., 2025). However, it deviates from other findings that suggested significant outshining (Sorba & Sawicki, 2015, 2018; Jain et al., 2024; Giménez-Arteaga et al., 2024; Narayanan et al., 2024). We attribute the primary systematic uncertainty in stellar mass estimation to the choice of SFH prior, rather than the resolved or integrated analysis.
A key direction for future works is the implementation of spatial binning strategies to mitigate the effects of PSF oversampling. While traditional binning approaches such as Voronoi tessellation (Cappellari & Copin, 2003) ensure a constant S/N, they do not fully exploit the spatial resolution of JWST imaging. We plan to incorporate the piXedfit binning methodology (Abdurro’uf & Akiyama, 2017; Abdurro’uf et al., 2021), which segments galaxies based on $\chi^{2}$ comparisons of neighbouring SEDs. This method retains detailed spatial information, while respecting PSF limitations and maintaining sufficient S/N within each bin, making it well suited to our framework.
In parallel, we aim to extend the pipeline to other JWST surveys, such as The Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al., 2025), and to explore datasets with different observational characteristics. Additional developments will include testing more flexible SFH priors and implementing pixel-level redshift inference. These efforts will enhance the robustness and interpretability of resolved stellar population studies in large galaxy samples.